OpenStudy (anonymous):

PLEASE EXPLAIN HOW YOU GOT THE ANSWER!!! Let P(x) be a polynomial such that P(5)=0. Which one of the following must be a factor of P(x)? A) x^2-25 B)x +25 C) x-5 D) x+5

OpenStudy (campbell_st):

well all you need to do is substitute x = 5 into each polynomial... the one that gives a zero answer is the solution to the question looking at D) x + 5 make the substitution 5 + 5 = 10 so D isn't the correct one hope it makes sense

OpenStudy (anonymous):

my lord you're a genius

OpenStudy (anonymous):

thanks man I owe you one

OpenStudy (anonymous):

wait actually hold on

OpenStudy (anonymous):

C) 5-5 =0 A) 25-25= 0

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

am I doing something wrong?

OpenStudy (campbell_st):

no you are correct... if you look at A x^2 - 25 = (x -5)(x + 5) and you know if P(x) = x-5 then P(5) = 0 so it seems there is a problem... I'd expect you'd need to contact the teacher... as for me there is a typo in the solutions... if they only want 1 answer